# Simple is Effective: The Roles of Graphs and Large Language Models in Knowledge-Graph-Based Retrieval-Augmented Generation

### v2

@@ B.1 Additional Implementation and Evaluation Details
$\textbf{z}_{\mathcal{N}(e)}^{l+1}=\textbf{MEAN}\left(\left\{[\textbf{z}_{e^{\prime}}^{l}\|\textbf{z}_{r}]\mid(e^{\prime},r,e)\in{\mathcal{G}}\right\}\right)$ (3)
$\textbf{z}_{e}^{l+1}=\sigma\left([\textbf{z}_{e}^{l}\|\textbf{z}_{\mathcal{N}(e)}^{l+1}\right)$ (4)
$\textbf{z}_{e}^{0}=\textbf{z}_{e}$ (5)
where $\sigma(\cdot)$ is an MLP. Empirically we find that 1-layer GraphSAGE yields the best performance.
B.2 Implementation Dependencies
@@
B.3 Additional Experiment Results for Varying K in Top-K Triple Retrieval
Figure 5: Retrieval effectiveness on WebQSP across a spectrum of $K$ values for top-$K$ triple retrieval.
@@ B.3 Additional Experiment Results for Varying K in Top-K Triple Retrieval
B.4 Evaluation Breakdown Over Topic Entity Count
@@
Model Shortest Path Triples GPT-4o-labeled Triples Answer Entities Single Multiple Single Multiple Single Multiple $(46.3\%)$ $(53.7\%)$ $(46.3\%)$ $(53.7\%)$ $(46.3\%)$ $(53.7\%)$ cosine similarity $0.586$ $0.403$ $0.587$ $0.499$ $0.517$ $0.638$ RoG $0.711$ $0.547$ $0.324$ $0.275$ $0.769$ $0.903$ G-Retriever $0.245$ $0.128$ $0.289$ $0.156$ $0.393$ $0.360$ MLP $0.788$ $0.568$ $0.777$ $0.600$ $0.894$ $0.873$ MLP + topic entity $0.858$ $0.690$ $0.819$ $0.718$ $0.894$ $0.889$ SubgraphRAG $\underline{\mathbf{0.877}}$ $\underline{\mathbf{0.754}}$ $\underline{\mathbf{0.871}}$ $\underline{\mathbf{0.820}}$ $\underline{\mathbf{0.911}}$ $\underline{\mathbf{0.916}}$
@@
Appendix C Supplementary KGQA Results
@@
To assess the impact of our retriever on overall performance as analyzed in Sec. 4, Table 9 provides additional ablation results using GPT-4o-mini as the reasoner.
Table 9: Ablation studies with different retrievers, using the same prompt and GPT-4o-mini as the reasoner. Rand refers to random triple sampling, RandNoAns removes triples with ground-truth answers after random sampling, and NoRetriever directly asks questions without KG info.
WebQSP CWQ WebQSP-sub CWQ-sub Macro-F1 Hit Macro-F1 Hit Macro-F1 Micro-F1 Hit Hit@1 Scoreh Macro-F1 Micro-F1 Hit Hit@1 Scoreh SubgraphRAG + Rand 47.70 68.37 33.13 39.82 47.15 23.74 68.57 64.59 69.67 35.55 34.26 42.94 40.48 54.81 SubgraphRAG + RandNoAns 36.83 49.63 25.69 30.70 35.77 14.62 48.94 44.96 57.12 26.25 26.11 31.60 29.21 48.80 SubgraphRAG + NoRetriever 47.49 71.01 33.43 42.25 46.68 25.53 70.94 62.67 60.22 35.66 31.11 44.66 39.99 44.24 SubgraphRAG + cosine similarity 64.94 78.19 41.23 49.22 65.26 43.84 78.90 74.15 73.03 45.07 42.27 53.83 49.37 57.06 SubgraphRAG + Retrieve-Rewrite-Answer 38.26 53.62 - - 37.54 19.62 53.37 50.10 67.68 - - - - - SubgraphRAG + StructGPT 71.62 82.68 - - 71.66 59.05 82.87 80.44 81.48 - - - - - SubgraphRAG + G-Retriever 59.28 75.25 36.79 42.23 59.55 35.19 76.01 72.87 76.31 39.55 39.42 45.47 43.29 57.67 SubgraphRAG + RoG-Sep 70.08 82.25 44.30 51.15 70.91 54.69 83.45 78.51 81.00 49.85 50.40 57.51 52.35 64.27 SubgraphRAG 77.45 90.11 54.13 62.02 78.34 58.44 91.34 87.36 82.21 61.13 58.86 70.01 65.48 64.20
Appendix D Prompt Templates
Fig. 6 is the detailed prompt template used in our experiments, and Fig. 7 provides the prompt employed to label relevant triples via GPT-4o.
@@ Appendix D Prompt Templates
User:   // ICL example Triplets: (Lou Seal,sports.mascot.team,San Francisco Giants) (San Francisco Giants,sports.sports_team.championships,2012 World Series) (San Francisco Giants,sports.sports_championship_event.champion,2014 World Series) (San Francisco Giants,time.participant.event,2014 Major League Baseball season) (San Francisco Giants,time.participant.event,2010 World Series) (San Francisco Giants,time.participant.event,2010 Major League Baseball season) (San Francisco Giants,sports.sports_team.championships,2014 World Series) (San Francisco Giants,sports.sports_team.team_mascot,Crazy Crab) (San Francisco Giants,sports.sports_team.championships,2010 World Series) (San Francisco Giants,sports.professional_sports_team.owner_s,Bill Neukom) (San Francisco Giants,time.participant.event,2012 World Series) (San Francisco,sports.sports_team_location.teams,San Francisco Giants) (San Francisco Giants,sports.sports_team.arena_stadium,AT&T Park) (AT&T Park,location.location.events,2012 World Series) (m.011zsc4_,organization.leadership.organization,San Francisco Giants) (San Francisco Giants,sports.sports_team.previously_known_as,New York Giants) (AT&T Park,location.location.events,2010 World Series) (Crazy Crab,sports.mascot.team,San Francisco Giants) (New York Giants,baseball.baseball_team.league,National League) (San Francisco Giants,sports.sports_team.colors,Black) (San Francisco Giants,sports.sports_team.previously_known_as,New York Gothams) (m.0k079qm,base.schemastaging.team_training_ground_relationship.team,San Francisco Giants) (m.0k079ry,base.schemastaging.team_training_ground_relationship.team,San Francisco Giants) (2010 World Series,time.event.locations,AT&T Park) (San Francisco Giants,time.participant.event,2012 Major League Baseball season) (San Francisco Giants,baseball.baseball_team.league,National League) (m.0crtd80,sports.sports_league_participation.league,National League West) (San Francisco Giants,sports.sports_team.location,San Francisco) (San Francisco Giants,sports.sports_team.sport,Baseball) (m.05n6dtn,baseball.baseball_team_stats.team,San Francisco Giants) Question: What year did the team with mascot named Lou Seal win the World Series?   Assistant:   // ICL example To find the year the team with mascot named Lou Seal won the World Series, we need to find the team with mascot named Lou Seal and then find the year they won the World Series. From the triplets, we can see that Lou Seal is the mascot of the San Francisco Giants. Now, we need to find the year the San Francisco Giants won the World Series. From the triplets, we can see that San Francisco Giants won the 2010 World Series and 2012 World Series and 2014 World Series. So, the team with mascot named Lou Seal (San Francisco Giants) won the World Series in 2010, 2012, and 2014. Therefore, the formatted answers are: ans: 2014 (2014 World Series) ans: 2012 (2012 World Series) ans: 2010 (2010 World Series)   User:  // Sample to infer Triplets: (ea, rab, eb)   // Retrieved triples (ec, rcd, ed) Question: what’s …?   // Q in datasets
Figure 6: Detailed prompt for LLM question-answering
@@
Based on the triplets retrieved from a knowledge graph, please select relevant triplets for answering the question. Please return formatted triplets as a list, each prefixed with “evidence:”.
@@
Appendix E Explainablity Examples
To demonstrate the superior explainability of our approach, we provide multiple example responses from our LLM reasoners below, which cover questions requiring different reasoning hops and logic chains. We also have included examples where our LLM reasoners refused to answer due to insufficient information, which helps to provide more truth-grounded answers.
@@ Appendix E Explainablity Examples
  LLM Response:
@@
  LLM Response:
@@
CWQ-WebQTest-1320_c5498ca807d2e1ec30d4c8fdd41f0bf7 Question:
What stadium does the team with mascot named Hank play at?
  LLM Response:
Miller Park
@@
  LLM Response:
@@
WebQSP-WebQTest-188 Question:
where did bristol palin go to school
  LLM Response:
West Anchorage High School
Wasilla High School Juneau-Douglas High School
WebQSP-WebQTest-760 Question:
what party does barack obama belong to
  LLM Response:
Democratic Party

### v3

@@ C.1 Additional Implementation and Evaluation Details
$\textbf{z}_{\mathcal{N}(e)}^{l+1}=\textbf{MEAN}\left(\left\{[\textbf{z}_{e^{\prime}}^{l}\|\textbf{z}_{r}]\mid(e^{\prime},r,e)\in{\mathcal{G}}\right\}\right),$ (4)
$\textbf{z}_{e}^{l+1}=\sigma\left([\textbf{z}_{e}^{l}\|\textbf{z}_{\mathcal{N}(e)}^{l+1}\right),$ (5)
$\textbf{z}_{e}^{0}=\textbf{z}_{e},$ (6)
where $\sigma(\cdot)$ is an MLP. Empirically, we find that 1-layer GraphSAGE yields the best performance.
C.2 Implementation Dependencies
@@
C.3 Additional Experiment Results for Varying K in Top-K Triple Retrieval
Figure 6: Retrieval effectiveness on WebQSP across various $K$ values for top-$K$ triple retrieval.
@@ C.3 Additional Experiment Results for Varying K in Top-K Triple Retrieval
C.4 Evaluation Breakdown Over Topic Entity Count
@@
Model Shortest Path Triples GPT-4o-labeled Triples Answer Entities Single Multiple Single Multiple Single Multiple $(46.3\%)$ $(53.7\%)$ $(46.3\%)$ $(53.7\%)$ $(46.3\%)$ $(53.7\%)$ cosine similarity $0.586$ $0.403$ $0.587$ $0.499$ $0.517$ $0.638$ RoG $0.711$ $0.547$ $0.324$ $0.275$ $0.769$ $0.903$ G-Retriever $0.245$ $0.128$ $0.289$ $0.156$ $0.393$ $0.360$ GNN-RAG $0.496$ $0.503$ $0.369$ $0.403$ $0.829$ $0.851$ MLP $0.788$ $0.568$ $0.777$ $0.600$ $0.894$ $0.873$ MLP + topic entity $0.858$ $0.690$ $0.819$ $0.718$ $0.894$ $0.889$ SubgraphRAG $\underline{\mathbf{0.877}}$ $\underline{\mathbf{0.754}}$ $\underline{\mathbf{0.871}}$ $\underline{\mathbf{0.820}}$ $\underline{\mathbf{0.911}}$ $\underline{\mathbf{0.916}}$
@@
(a) WebQSP-sub
(b) CWQ-sub
Figure 7: Ablation studies on the number of retrieved triples and the used retrievers in SubgraphRAG. The triples from RoG’s retriever are generated using beam search with sizes {3, 6, 9, 12}, and all the experiments use SubgraphRAG’s un-fine-tuned LLM reasoners.
Appendix D Supplementary KGQA Results
@@
Impact of Retrieval Size. Fig. 4 shows how our SubgraphRAG would perform with different numbers of retrieved triples. To further demonstrate the superior performance of our proposed retriever module, we also scale RoG’s retriever with more retrieved triples by increasing its beam search sizes (noting that this substantially increases computational costs). As shown in Fig. 7, clearly, even when using RoG’s retriever with more triples in SubgraphRAG, our proposed retriever consistently achieves significantly better results, further demonstrating the effectiveness of our retriever module.
Table 9: Ablation studies with different retrievers, using the same prompt and GPT4o-mini as the reasoner. Rand refers to random triple sampling, RandNoAns removes triples with ground-truth answers after random sampling, and NoRetriever directly asks questions without KG info.
WebQSP CWQ WebQSP-sub CWQ-sub Macro-F1 Hit Macro-F1 Hit Macro-F1 Micro-F1 Hit Hit@1 Scoreh Macro-F1 Micro-F1 Hit Hit@1 Scoreh Rand + SubgraphRAG Reasoner 47.70 68.37 33.13 39.82 47.15 23.74 68.57 64.59 69.67 35.55 34.26 42.94 40.48 54.81 RandNoAns + SubgraphRAG Reasoner 36.83 49.63 25.69 30.70 35.77 14.62 48.94 44.96 57.12 26.25 26.11 31.60 29.21 48.80 NoRetriever + SubgraphRAG Reasoner 47.49 71.01 33.43 42.25 46.68 25.53 70.94 62.67 60.22 35.66 31.11 44.66 39.99 44.24 cosine similarity + SubgraphRAG Reasoner 64.94 78.19 41.23 49.22 65.26 43.84 78.90 74.15 73.03 45.07 42.27 53.83 49.37 57.06 Retrieve-Rewrite-Answer + SubgraphRAG Reasoner 38.26 53.62 - - 37.54 19.62 53.37 50.10 67.68 - - - - - StructGPT + SubgraphRAG Reasoner 71.62 82.68 - - 71.66 59.05 82.87 80.44 81.48 - - - - - G-Retriever + SubgraphRAG Reasoner 59.28 75.25 36.79 42.23 59.55 35.19 76.01 72.87 76.31 39.55 39.42 45.47 43.29 57.67 RoG-Sep + SubgraphRAG Reasoner 70.08 82.25 44.30 51.15 70.91 54.69 83.45 78.51 81.00 49.85 50.40 57.51 52.35 64.27 GNN-RAG + SubgraphRAG Reasoner 73.27 85.93 52.91 59.64 74.24 53.92 87.24 83.39 84.41 60.89 59.86 68.43 64.04 67.71 GNN-RAG ($\leftrightarrow$) + SubgraphRAG Reasoner 61.19 78.99 43.42 51.37 61.41 42.02 79.79 74.98 76.11 48.04 43.7 56.88 52.32 59.45 SubgraphRAG 77.45 90.11 54.13 62.02 78.34 58.44 91.34 87.36 82.21 61.13 58.86 70.01 65.48 64.20 SubgraphRAG ($\leftrightarrow$) 73.81 88.08 44.69 54.21 74.42 49.41 89.10 84.67 81.35 49.47 45.16 60.18 54.18 58.86
Appendix E Prompt Templates
Fig. 8 is the detailed prompt template used in our experiments for all samples, and Fig. 9 provides the prompt employed to label relevant triples via GPT-4o.
@@ Appendix E Prompt Templates
User:   // ICL example Triplets: (Lou Seal,sports.mascot.team,San Francisco Giants) (San Francisco Giants,sports.sports_team.championships,2012 World Series) (San Francisco Giants,sports.sports_championship_event.champion,2014 World Series) (San Francisco Giants,time.participant.event,2014 Major League Baseball season) (San Francisco Giants,time.participant.event,2010 World Series) (San Francisco Giants,time.participant.event,2010 Major League Baseball season) (San Francisco Giants,sports.sports_team.championships,2014 World Series) (San Francisco Giants,sports.sports_team.team_mascot,Crazy Crab) (San Francisco Giants,sports.sports_team.championships,2010 World Series) (San Francisco Giants,sports.professional_sports_team.owner_s,Bill Neukom) (San Francisco Giants,time.participant.event,2012 World Series) (San Francisco,sports.sports_team_location.teams,San Francisco Giants) (San Francisco Giants,sports.sports_team.arena_stadium,AT&T Park) (AT&T Park,location.location.events,2012 World Series) Question: What year did the team with mascot named Lou Seal win the World Series?   Assistant:   // ICL example To find the year the team with mascot named Lou Seal won the World Series, we need to find the team with mascot named Lou Seal and then find the year they won the World Series. From the triplets, we can see that Lou Seal is the mascot of the San Francisco Giants. Now, we need to find the year the San Francisco Giants won the World Series. From the triplets, we can see that San Francisco Giants won the 2010 World Series and 2012 World Series and 2014 World Series. So, the team with mascot named Lou Seal (San Francisco Giants) won the World Series in 2010, 2012, and 2014. Therefore, the formatted answers are: ans: 2014 (2014 World Series) ans: 2012 (2012 World Series) ans: 2010 (2010 World Series)   User:  // Sample to infer Triplets: (ea, rab, eb)   // Retrieved triples (ec, rcd, ed) Question: what’s …?   // Q in datasets
Figure 8: Detailed prompt for LLM question-answering used in our experiments for all samples.
@@
Based on the triplets retrieved from a knowledge graph, please select relevant triplets for answering the question. Please return formatted triplets as a list, each prefixed with ”evidence:”.
@@
Appendix F Explainablity Examples
To demonstrate the superior explainability of our approach, we provide multiple example responses from our LLM reasoners below, which cover questions requiring different reasoning hops and logic chains. We have also included examples of our LLM reasoners refusing to answer due to insufficient information in Appendix F.2, which helps to provide more truth-grounded answers. Due to limited space, we do not include retrieved triples for all the examples, but one will be able to find full prediction files in our released repository.
F.1 LLM Responses with Explainable Reasoning
@@
  Retrieved Triples:
(Haiti,location.country.languages_spoken,Haitian Creole)
(Haiti,location.country.first_level_divisions,Nord-Ouest Department) (Haiti,location.country.official_language,French) (French,language.human_language.countries_spoken_in,Haiti) (Nord-Ouest Department,location.administrative_division.first_level_division_of,Haiti) (Haiti,location.country.languages_spoken,French) (Haiti,location.country.administrative_divisions,Nord-Ouest Department) (Haiti,location.country.official_language,Haitian Creole) (Nord-Ouest Department,location.location.containedby,Haiti) (Haitian Creole,language.human_language.main_country,Haiti)   LLM Response:
@@
CWQ-WebQTest-1320_c5498ca807d2e1ec30d4c8fdd41f0bf7 Question:
What stadium does the team with mascot named Hank play at?
  Retrieved Triples:
(Milwaukee Brewers,sports.sports_team.team_mascot,Hank)
(Hank,sports.mascot.team,Milwaukee Brewers) (Milwaukee Brewers,sports.sports_team.previously_known_as,Seattle Pilots) (Seattle Pilots,sports.defunct_sports_team.later_known_as,Milwaukee Brewers) (Milwaukee Brewers,sports.sports_team.arena_stadium,Miller Park) (Milwaukee Brewers,sports.sports_team.location,Milwaukee) (Miller Park,sports.sports_facility.teams,Milwaukee Brewers) (Milwaukee,sports.sports_team_location.teams,Milwaukee Brewers) (Miller Park,location.location.containedby,Milwaukee) (Milwaukee Brewers,baseball.baseball_team.current_coaches,m.0527c88)   LLM Response:
Miller Park
@@
  Retrieved Triples:
(Texas Rangers,sports.professional_sports_team.owner_s,Tom Hicks)
(Tom Hicks,sports.sports_team_owner.teams_owned,Texas Rangers) (Dallas Stars,sports.professional_sports_team.owner_s,Tom Hicks) (Seattle Pilots,sports.defunct_sports_team.later_known_as,Milwaukee Brewers) (Tom Hicks,sports.sports_team_owner.teams_owned,Dallas Stars) (United States of America,base.locations.countries.states_provinces_within,Alabama) (Houston Astros,sports.professional_sports_team.owner_s,Drayton McLane, Jr.) (m.0crtd2r,sports.sports_league_participation.team,Los Angeles Angels of Anaheim) (United States of America,base.aareas.schema.administrative_area.administrative_children,Alabama) (m.0crtd31,sports.sports_league_participation.team,Anaheim Angels) (United States of America,base.aareas.schema.administrative_area.administrative_children,Washington, D.C.) (United States of America,base.locations.countries.states_provinces_within,Mississippi) (m.0crtd3v,sports.sports_league_participation.team,California Angels) (United States of America,base.aareas.schema.administrative_area.administrative_children,Mississippi) (American League West,baseball.baseball_division.teams,Texas Rangers)   LLM Response:
@@
  Retrieved Triples:
(Sweden,location.location.time_zones,Central European Time Zone)
(Sweden,location.country.languages_spoken,Finnish Language) (Finnish Language,language.human_language.countries_spoken_in,Sweden) (Sweden,location.country.languages_spoken,Turoyo Language) (Sweden,location.country.official_language,Swedish Language) (Spain,location.location.time_zones,Central European Time Zone) (Stockholm,location.location.time_zones,Central European Time Zone) (Unitary state,government.form_of_government.countries,Sweden) (Swedish Language,language.human_language.countries_spoken_in,Finland) (Sweden,location.country.languages_spoken,Swedish Language)   LLM Response:
@@
WebQSP-WebQTest-188 Question:
where did bristol palin go to school
  Retrieved Triples:
(m.0byl4zl,education.education.institution,Wasilla High School)
(Wasilla High School,education.educational_institution.students_graduates,m.0byl4zl) (Juneau-Douglas High School,education.educational_institution.students_graduates,m.0cpkbwp) (m.0cpkbwp,education.education.institution,Juneau-Douglas High School) (Bristol Palin,people.person.education,m.0byl4zl) (Bristol Palin,people.person.education,m.0cpk4bl) (Bristol Palin,people.person.education,m.0cpkbwp) (m.0cpk4bl,education.education.institution,West Anchorage High School) (West Anchorage High School,education.educational_institution.students_graduates,m.0cpk4bl) (m.0byl4zl,education.education.student,Bristol Palin) (m.0cpk4bl,education.education.student,Bristol Palin) (m.0cpkbwp,education.education.student,Bristol Palin) (m.04tlm1z,education.education.institution,Wasilla High School) (m.0n1g77g,education.education.institution,Wasilla High School) (Wasilla High School,education.educational_institution.students_graduates,m.04tlm1z) (Wasilla High School,education.educational_institution.students_graduates,m.0n1g77g) (Bristol Palin,people.person.place_of_birth,Wasilla) (Wasilla,location.location.people_born_here,Bristol Palin) (Bristol Palin,people.person.nationality,United States of America) (Levi Johnston,people.person.place_of_birth,Wasilla)   LLM Response:
West Anchorage High School
Wasilla High School Juneau-Douglas High School
WebQSP-WebQTest-760 Question:
what party does barack obama belong to
  Retrieved Triples:
(Barack Obama,influence.influence_node.influenced_by,Saul Alinsky)
(Barack Obama,organization.organization_founder.organizations_founded,National Commission on Fiscal Responsibility and Reform) (Barack Obama presidential campaign, 2012,government.election_campaign.party,Democratic Party) (Barack Obama,government.politician.election_campaigns,Barack Obama presidential primary campaign, 2008) (Barack Obama,government.politician.party,m.03gjhww) (Barack Obama,government.politician.election_campaigns,Barack Obama 1996 Illinois Senate Campaign) (Barack Obama 2004 US Senate Campaign,government.election_campaign.party,Democratic Party) (Barack Obama,influence.influence_node.influenced_by,Reinhold Niebuhr) (m.03gjhww,government.political_party_tenure.party,Democratic Party) (Barack Obama,organization.organization_founder.organizations_founded,White House Task Force to Protect Students from Sexual Assault) (Barack Obama presidential primary campaign, 2008,government.election_campaign.party,Democratic Party) (Barack Obama 1996 Illinois Senate Campaign,government.election_campaign.party,Democratic Party) (Barack Obama,base.politicalconventions.presidential_nominee.nominated_at,2008 Democratic National Convention) (Barack Obama,government.politician.government_positions_held,m.03h96h6) (Barack Obama,government.politician.election_campaigns,Barack Obama 2004 US Senate Campaign)   LLM Response:
Democratic Party
@@
F.2 LLM Refusal to Answer Due to Insufficient Information